jdvaral:

152x-378y=-74 and -378x+152y=-604 what is x+y:x-y=?

jhonyy9:

like a first step try factorizing out what is posibile 152x -378y = -74 2*76x -2*189y = -2*37 2(76x-189y) = -2*37 divide both sides by 2 76x -189y = -37 with second equation make the same -378x +152y =-604 -2*189x +2*76y = -2*302 divide both sides by 2 -189x +76y = -302 and now there are 76x -189y =-37 -189x +76y = -302 @jdvaral please rewrite the question ! this is (x+y)/(x-y) = ?